OpenStudy (anonymous):

Which is the best buy? A) A) 5 oz for 42¢ B) B) 24 oz for $1.68 C) C) 18 oz for $1.17 D) D) 26 oz for $2.39

OpenStudy (anonymous):

You have to determine how much costs each onze for each option, for example, in option A \[5oz=42 c\] you divide by 5 \[1 oz=\frac{ 42 }{ 5 } c=8.4c\] now, do the same for B, C and D

OpenStudy (jdoe0001):

well, the better "ratio" per pennie will be 5/42 = 0.12 24/168 = 0.143 18/117 = 0.154 26/239 = 0.11

OpenStudy (jdoe0001):

so I'd say you get more "ounces per pennies" on the 18ouncer

OpenStudy (anonymous):

well I agree with jdoe

OpenStudy (anonymous):

Thanks!

OpenStudy (jdoe0001):

yw